OpenStudy (anonymous):

Can the expression s times the square root of the quantity 64 times s cubed. + 2stthe square root of s. + 8the square root of the quantity 2 times s. be simplified into an expression with fewer terms? Explain and simplify, if possible.

OpenStudy (anonymous):

@Jhannybean can you help me?

OpenStudy (jhannybean):

\[\large s\times s^3\times \sqrt{64}\] I don't understand the second part. "+ 2stthe square root of s. + 8the square root of the quantity 2 times s. "

OpenStudy (anonymous):

Let me draw it for you....

OpenStudy (anonymous):